# Phase transition like behaviors of Propagation of Passenger Stranding phenomena in Subway Networks

### 12 Jan 2024 v1

@@ Abstract
The subway system as the most important transportation for daily urban commuting is a typical non-equilibrium complex system. Unlike complex systems composed of a single type of basic units, the subway system consists of 2 types of basic units with service relationship: passengers as the demander of services and trains as the supplier of services. The main challenge of the operation of subway system arises from passenger stranding at stations, which arise from the spatiotemporal mismatch of supply scale and demand scale. There is a special phenomenon of the propagation of passenger stranding (PPS) within the stations in the system by forming stranded stations clusters under the condition of shortage of supply, which significantly reduces the efficiency of the subway system. In this study, we take the Beijing subway as an example to study the nature of PPS phenomena from a view point of statistical physics, in which both an ideal model and a realistic data-driven model are considered. The simulation results suggest that both the efficiency of subway service (as measured by passenger waiting time) and the scale of stranding demonstrate phase transition like behaviors. The phase transition point can quantitatively characterize the resilience failure threshold of subway service. PPS phenomenon encompasses a multitude of coupling effects within the subway system. The eigen microstate method is used to extracting the fundamental patterns of the formation and regulation mechanism of stranded stations clusters. The research in this study provides a theoretical basis for understanding the mechanism of service efficiency in subway networks and offers a theoretical foundation for strategies to improve service, such as topological planning and train timetable optimization. Moreover, the methodology used in present work has significant implications for service systems.
Keywords First keyword  $\cdot$ Second keyword  $\cdot$ Third keyword More
@@ 1 Introduction
Among various urban transportation modes, the subway’s travel time is the most easily controlled due to its immunity from issues like road congestion and inclement weather. The carbon emissions per passenger transported and unit distance covered by electric-powered subways are significantly lower than other urban transportation modes. As a result, subways offer an optimal solution to address common urban issues such as traffic congestion, environmental pollution, and limited efficiency of public services in large cities. The spatial extension characteristics of subways also shape the structure of large cities and the characteristics of residents’ activities, known as Transit-Oriented Development (TOD)[1] . By utilizing the economy and efficient subway travel mode, residents can opt to reside in low-cost suburbs while engaging in employment in the city center, resulting in a significant spatial decoupling between job and residence centers. As the most crucial public service system in cities, it is required that the subway system should provide efficient services to meet the highly concentrated daily commuting demands. The efficiency of a subway system is primarily characterized by its ability to fulfill the transportation demands of passengers in a timely manner, as well as minimize passenger waiting time.
@@
Because the arrival of trains as service providers is not continuous but discrete in time, and the arrival of passengers as service demanders at the station is random, there are always random events of passenger stranded at the platform, and there is no temporal correlation between these events at stations. When the passenger arrival rate at a station exceeds the efficiency of subway’s service, the events of passenger stranding start to exhibit temporal correlation. During the morning rush hours, the Beijing subway system encounters the most severe instances of stranding, with 8.547% of stations affected (Fig.1 A illustrates the number of stations with stranding phenomena over time on Monday, May 10, 2021). The stranding phenomenon can propagate from station to station along the subway line, and these stranded stations forming a spatially correlated cluster. We call this phenomenon as the propagation of stranded passengers (PPS). PPS comprises cluster formation. As an example, the stranded stations during the morning rush hours are illustrated in Fig.1 B. The presence of temporal and spatial correlations of stranding in the subway not only leads to longer commuting times for passengers, but also undermines the advantages of subway travel and depresses societal efficiency. When the duration and spatial scale of stranding reach a substantial magnitude, the system is unable to provide adequate service supply to meet the demand. The demand scale in this case exceeds the threshold of system resilience. Large-scale passenger stranding within enclosed subway stations increases the difficulty of evacuation and raises concerns regarding public safety, such as the spread of epidemics and the risk of terrorist attacks.
@@
The critical demand scale where the subway system fails to provide services is defined as the threshold of service resilience, which reflects the ability to meet demand and serves as a crucial parameter for quantitative evaluation in decision-making processes related to investment, insurance, risk management, and emergency preparedness. The term "resilience" originated in mechanics to describe the resistance of materials to physical impacts. In the study of complex systems, the concept of resilience can be primarily categorized into three frameworks[2] . (1) System resilience (or ecological resilience) focuses on the magnitude of the change or perturbation that a system can endure without shifting to another stable state[3] . (2) Engineering resilience is defined by the recovery rate or time[4]. (3) Adaptive resilience characterizes the capacity of socioecological systems to adapt or transform in response to unfamiliar, unexpected and extreme shocks[5] . The resilience of services can be attributed to the category of system resilience. In present work, the system’s states are characterized by its responsive properties, with demand scale serving as the stimulus and the scale of stranding representing the system’s response. The threshold of resilience is defined as the demand scale at which a significant change in response behavior occurs, at which the system can no longer provide normal service beyond a certain threshold.
The subway system is a non-equilibrium complex system composed of a vast number of interacting basic units (passengers and trains), making it a typical subject of study in statistical physics[6] . Unlike the systems composed of a single type of unit, the subway system is a service system consisting of two distinct types of units interplay through service relationship: (1) passengers as the demander of services and (2) trains as the supplier of services. A demand of service from a single passenger is a pair of stations depending on time called as origin-destination (OD) pairs, and on the supply side of services, subway trains operate according to the schedules of different lines. Stranding phenomenon of passengers at stations and the propagation of stranding within the network are the result of collective behaviors of both passengers and trains, which are from the alignment of supply and demand at each station in a timely manner, as well as the spatial correlation between stations facilitated by trains. For the condition of supply scarcity, the stranded stations become interconnected to form clusters. The generation of these clusters relies on a nucleation and growth like processes[7] , which is similar with the condensation of gas molecules attracting and coalescing into a liquid state during vapor-liquid transitions, and the growth process of magnetization in ferromagnetic materials, exhibit similarities to phase transition phenomena. To reveal the mechanism underlying the PPS phenomenon, it is essential to conduct research employing statistical physics theories and methodologies.
@@
Generally, the subway was studied by traffic fluid model[8] or complex networks[8]. The traffic fluid model is inspired by the studies of congestion in road traffic[9; 10; 11; 12; 13; 14]. The fundamental diagram of traffic flow and velocity reflects the spontaneous phase transition of congestion[15]. Subway studies based on complex networks investigate the reliability of efficient travel paths by establishing undirected network models[16; 17] or directed network models[18], considering the network connectivity characteristics of stations. Unlike traffic flow in roads and transportation problems in complex networks, congestion phenomena in subway systems arise from the spatiotemporal matching of service and demand. Insufficient service provision results in passengers being stranded. The relaxation time of stranding phenomena at individual stations is governed by the ratio of local service supply rate and the demand rate. The service supply rate is determined by the passenger capacity of trains and the arrival rate of trains (timetable), and the demand rate is determined by the arrival rate of passengers at this station. The development patterns of TOD lead to highly spatiotemporally uneven demands for subway services due to the tidal characteristics of commuting during morning and evening rush hours[19; 20]. This makes it challenging to achieve spatial and temporal matching between service supply and demand. Besides the local mismatching of demand and supply, it is worth noting that the stranding conditions at upstream stations would reduce the supply rate of downstream station as well due to the characteristics of the train providing service along the lines, as depicted in Fig.1 D. When the train operating at almost full capacity, the number of passengers getting off the train might lower than that who want to get on which gives rise to the propagation of stranding phenomena along the line. The TOD mode contributes to a significant spatial separation between job and residence centers, resulting in a large average commuting distance for passengers which determines the spatial correlation length of such stranding phenomena.
In order to determine the resilience threshold of a system, it is necessary to investigate the response behavior of subway systems under varying demand scales. However, executing controlled variable experiments within the actual infrastructure of mega-cities, particularly when aiming to surpass the resilience threshold, encounters substantial obstacles. As a typical non-equilibrium system, a data-driven dynamic model can be established, enabling the acquisition of its response patterns under different demand scales through computer simulations. In turn, the service resilience threshold can be predicted. In present work, the system’s responsive behaviors to varying demand scales are investigated in both steady-state and non-steady-state non-equilibrium conditions. The responsive behaviors are characterized by both the scale of stranding within the system and the average waiting time of passengers. The steady-state conditions are investigated under the assumption of spatial and temporal homogeneous in both demand and supply. In this condition, the supply rate is fixed and the average demand rate (Pu) at each station is varied to reveal the general nature of the PPS phenomenon arising from supply shortage. Additionally, to simulate non-steady-state conditions, we incorporate time-dependent demand and supply rates with rush hour features in both the morning and evening, obtained from real-system data. Similar to the steady-state conditions, the supply rate is fixed as well and the demand rate, Pr is varied. Here, Pr is defined by the ratio of simulated passenger arrival rate to actual passenger arrival rates.
@@
PPS phenomenon encompasses a multitude of coupling effects within the subway system. Extracting the fundamental patterns from these coupling effects can provide valuable insights for the development of subway infrastructure, regulation of social services, and the establishment of effective methods for emergency response and incident prediction. In the field of non-equilibrium statistical physics, eigen microstate method offers a powerful method to investigate phase transitions, critical phenomena, and dynamic evolutions in complex systems, thereby providing a new avenue for studying PPS phenomena. By applying the eigen microstates method[21; 22; 23; 24; 25; 26], fundamental patterns and their evolutionary rules can be extracted from extensive system evolution data. In present work, PPS as a class of emergent behaviors in self-organized multi-body systems are studied from a statistical physics perspective, using eigen microstates method to uncover the underlying mechanisms.
@@
Service system can be considered as a general model for many economic and social activities of human beings. In these systems the scheduling of supply is optimized to meet the spatiotemporal characteristics of demand and PPS phenomena is a common feature of these system. Furthermore, the global operational efficiency, cost-effectiveness and prevent system from collapse are essential considerations from society management point of view. Examples of service systems include transportation systems, like subway, taxis[27] , buses[28; 29] etc; elevators in skyscrapers[30; 31; 32; 33]; medical appointment systems; logistical management[34; 35]; aircraft carriers decks management[36; 37; 38]. In present work, the subway system as a reprehensive example of service system is studied and the mechanism of PPS phenomena of service system is revealed. The results and methodologies may provide valuable insights for service systems.
@@
Figure 1: Stranding phenomena is a major problem of subway system: (A) The temporal variation in the number of stations with stranded passengers in the subway network throughout a day on May 10, 2021; (B) Spatial distribution of stations with stranded passengers in the Beijing subway network at 8:10 AM rush hour; (C) Mechanisms of passengers being stranded at subway stations. The causes for the occurrence of stranded passengers and PPS phenomena are as follows: network topology, service trains scheduling, and spatiotemporal characteristics of demand. (D) the scheme of causing of PPS phenomena. stranding occurs at stations when passengers’ demand exceeds transportation supply. Due to the interconnected nature of the line, operation in full capacity in upstream stations would occupy the supply scale of downstream stations, leading to passengers stranded at the latter.
2 Result
@@
To investigate the phenomenon of PPS in subway systems, we have developed a dynamic model based on an actual subway network, using Beijing as a case study. Beijing, a world-class megacity with a population of 23 million permanent residents. As of July 26, 2022, the Beijing subway network composes of 17 lines with a total of 330 operational stations and 69 transfer stations, constituting a transportation system transportation system of significant scale. The job and residence centers in Beijing exhibit characteristic features of TOD, with a substantial portion of the population relying on the subway for their daily commuting demands. In present work, a model system includes 3 east-west linear lines (Lines 1, 6, and 7), 3 north-south linear lines (Lines 4, 5, and 8), and 3 circular lines (Lines 2, 10, and 13) are selected, resulting in a subway network comprising 234 stations (N=234). Among these stations, there are 36 transfer stations and 10 terminal stations (ends), capturing the representative topological structure of the Beijing subway system. These 9 earliest established lines, operating in the core area of Beijing serve as the primary main arteries in the real subway system. This model system is an appropriate representation in terms of scale, significance, and topological structure, which enables us to capture the daily operational features. In order to concentrate on the primary physical properties and alleviate computational burden, we have excluded the newly constructed lines from our study that have not yet achieved supply-demand equilibrium.
@@
Within this subway model, a dynamic evolutionary model system to quantitatively investigate the phenomenon of PPS is put forward. This model system comprises two types of active basic units, passengers ($N_{p}=3,811,418$) and trains ($N_{t}=5,266$) with service relationship. The service process includes the operation of trains to meet the demand of passengers. Trains operate according to a predetermined schedule, ensuring punctuality. When a train arrives at a station, passengers on board make decisions to disembark or not according to their destination demand, while passengers waiting at the station Passengers decide whether to get onboard based on the crowdedness of the compartment. Through computer simulations, the dynamic trajectories of passengers and trains can be obtained (further details can be found in Supplement Meterials, A.2). To elucidate the fundamental physical characteristics of PPS, 2 simulation systems are considered: an ideal model (System I) and a realistic model (System II). In System I, it is assumed that the rates of demand and supply are independent of time and position, and the positions of OD pairs are randomly distributed throughout the network. This model system allows us to study PPS under non-equilibrium steady-state conditions. In System II, the rate of passenger demand depending time exhibiting with rush hour patterns. The positions of OD pairs capture the tidal nature of real commuting scenarios. The simulation time is measured in seconds, with the system state recorded every 5 minutes for System I and every 10 minutes for System II.
@@
This study employs the average waiting time of passengers and the scale of stranding as the order parameters characterizing the PPS phenomena. The waiting time of passengers, $T_{w}$, will increase due to the PPS phenomena, which serves as a crucial metric for evaluating the system’s service efficiency. To quantify the PPS phenomena in the subway network, we develop here a quantitative framework based on a physical concept of cluster, which combines evolving traffic dynamics with network structure. Instead of the commonly used structural topology, only stations in the network with stranded passengers are considered functionally connected. In this way, we can characterize and understand the dynamics of PPS phenomena through cluster formation and vanishing. The scale of stranding in the system is characterized by 3 order parameters: the total number of stations within the stranding clusters, n_stranded; the size of the largest cluster, max_cluster, represented by the number of stations within it; and the total number of stranded stations in the system, tot_stranded(quantitative definition of clusters can be found in Supplement Meterials, B.1.

### 28 Feb 2024 v2

@@ Abstract
The subway as the most important transportation for daily urban commuting is a typical non-equilibrium complex system, composed of 2 types of basic units with service relationship. One challenge to operate it is passengers be stranded at stations, which arise from the spatiotemporal mismatch of supply scale and demand scale. More seriously, there is a special phenomenon of the propagation of passenger stranding (PPS) by forming stranded stations clusters, which significantly reduces the service efficiency. In this study, Beijing subway as an example is studied to reveal the nature of PPS phenomena from a view point of statistical physics. The simulation results demonstrate phase transition like behaviors depending on the ratio of service supply scale and demand scale. The transition point can quantitatively characterize the resilience failure threshold of service. The eigen microstate method is used to extracting the fundamental patterns of PPS phenomena. Moreover, this study offers a theoretical foundation for strategies to improve service, such as topological planning and train timetable optimization. The methodology developed in present work has significant implications for study of other service systems.
@@ 1 Introduction
Among various urban transportation modes, the travelling time by taking the subway is the most easily controlled due to its immunity from issues like road congestion and inclement weather. The carbon emissions per passenger transported and unit distance covered by electric-powered subways are significantly lower than other options. As a result, subways offer an optimal solution to address common urban issues such as traffic congestion, environmental pollution, and limited efficiency of public services in large cities. The spatial extension characteristics of subways also shape the structure of large cities and the characteristics of residents’ activities, known as Transit-Oriented Development (TOD)1. By taking advantage of the economy and efficiency of the subway, residents can opt to reside in low-cost suburbs while pursuing employment opportunities in the city center, resulting in a significant spatial decoupling between job and residence centers. As the most crucial public service system in cities, the subway system bears the responsibility to provide efficient services to meet the highly concentrated daily commuting demands. Its efficiency is primarily characterized by its ability to fulfill the transportation demands of passengers while effectively reducing waiting time.
@@
Because of the discrete arrival of trains as service providers, coupled with the randomness of passenger arrivals as service demanders, there are invariably random incidents of passengers being stranded in stations. Generally, there is no temporal correlation between these events at stations. When the passenger arrival rate at a station exceeds the efficiency of the subway’s service, the events of passenger stranding start to exhibit temporal correlation. During the morning rush hours, the Beijing subway system encounters the most severe instances of stranding, with 8.547% of stations affected. Fig. 1a illustrates the number of stations with stranding phenomena over time on Monday, May 10, 2021. The stranding phenomenon can propagate from station to station along the subway line, and these stranded stations form a spatially correlated cluster. We call this phenomenon the propagation of stranded passengers (PPS). As an example, the stranded stations during the morning rush hours are illustrated in Fig. 1b. The presence of temporal and spatial correlations of stranding in the subway not only leads to longer commuting times for passengers, but also undermines the advantages of subway travel and depresses societal efficiency. Large-scale passenger stranding within enclosed subway stations increases the difficulty of evacuation and raises concerns regarding public safety, such as the spread of epidemics and the risk of terrorist attacks.
@@
When the duration and spatial scale of stranding reach a substantial magnitude, the system is unable to provide adequate service supply to meet the demand. The critical demand scale in this case can be defined as the threshold of service resilience, which serves as a crucial parameter for quantitative evaluation in decision-making processes related to investment, insurance, risk management, and emergency preparedness. The term “resilience” originated in mechanics to describe the resistance of materials to physical impacts. In the study of complex systems, the concept of resilience can be primarily categorized into 3 frameworks2: (1) System resilience (or ecological resilience) focuses on the magnitude of the change or perturbation that a system can endure without shifting to another stable state3; (2) Engineering resilience is defined by the recovery rate or time4; (3) Adaptive resilience characterizes the capacity of socioecological systems to adapt or transform in response to unfamiliar, unexpected and extreme shocks5. The resilience of services studied here can be attributed to the category of system resilience. The system’s responsive properties, as characterized by the stranding scale, are measured by adjusting the demand scale, which acts as the stimulus. The threshold of resilience is defined as the demand scale at which a notable change in response behavior occurs, beyond which the service efficiency of the system is significantly decreased.
@@
Generally, the subway has been studied using traffic fluid models6 or complex networks6. The traffic fluid models are inspired by the studies of congestion in road traffic7, 8, 9, 10, 11, 12. The fundamental diagram of traffic flow and velocity reflects the spontaneous phase transition of congestion5. Subway studies based on complex networks investigate the reliability of efficient travel paths by establishing undirected network models13, 14 or directed network models15, considering the network connectivity characteristics of stations. Unlike traffic flow in roads and transportation problems in complex networks, the PPS phenomenon stem from the timely fulfillment of service relationships locally at each station governed by the ratio of service supply rate and the demand rate, as well as the factors of topological of network as shown in Fig. 1c. The service supply rate is determined by the passenger capacity of trains and the arrival rate of trains (timetable), and the demand rate is determined by the arrival rate of passengers at this station. The development patterns of TOD lead to highly spatiotemporally uneven demands for subway services due to the tidal characteristics of commuting during morning and evening rush hours16, 17. This makes it challenging to achieve spatial and temporal matching between service supply and demand. The topological factor depends on both the connections of the stations to the network and the lines along which the trains provide service. This is because the passengers cannot travel freely within the networks, but must rely on trains to travel along predefined lines. If the planned route of a demand needs to pass through 2 or more subway lines, transferring to different trains is needed. Besides the local mismatching of demand and supply, it is worth noting that the stranding conditions at upstream stations would reduce the supply rate of downstream station as well due to the characteristics of the train providing service along the lines, as depicted in Fig. 1d. When the train operates at almost full capacity, the number of passengers getting off the train might be lower than those who want to get on, which gives rise to the propagation of stranding phenomena along the line. The TOD mode contributes to a significant spatial separation between job and residence centers, resulting in a large average commuting distance for passengers, which determines the spatial correlation length of PPS phenomena.
The transportation in subway system is a collective behavior of a vast number of active basic units (passengers and trains), making it a typical subject of study in on-equilibrium statistical physics18. Different from the systems composed of a single type of unit, the subway system is a service system consisting of 2 distinct types of units interplay through service relationship: (1) passengers as the demander of services and (2) trains as the supplier of services. A demand of service from a single passenger is a pair of stations depending on time called as origin-destination (OD) pairs, and on the supply side of services, subway trains operate according to the schedules of different lines. The formation of clusters of stranded stations, which serves as the main characteristic of PPS19, exhibits similarities to phase transition phenomena similar to the condensation of gas molecules attracting and coalescing into a liquid state during vapor-liquid transitions, and the growth process of magnetization in ferromagnetic materials. To reveal the mechanism underlying the PPS phenomenon, it is essential to conduct research employing statistical physics theories and methodologies.
@@
In order to determine the service resilience threshold, it is necessary to investigate the response behavior of subway systems under varying demand scales. However, executing controlled variable experiments within the actual infrastructure of large cities, particularly when aiming to surpass the resilience threshold, encounters substantial obstacles. As a typical non-equilibrium system, a data-driven dynamic model can be established, enabling the acquisition of its response patterns under different demand scales through computer simulations. In turn, the service resilience threshold can be predicted. In the present work, the system’s responsive behaviors which are characterized by both the scale of stranding and the average waiting time of passengers are investigated in both steady-state and non-steady-state non-equilibrium conditions. Moreover, the PPS phenomenon encompasses a multitude of coupling effects within the subway system. Extracting the fundamental patterns from these coupling effects can provide valuable insights for the development of subway infrastructure, regulation of social services, and the establishment of effective methods for emergency response and incident prediction. In the field of non-equilibrium statistical physics, eigen microstate method offers a powerful method to investigate phase transitions, critical phenomena, and dynamic evolutions in complex systems, thereby providing a powerful tool for studying PPS phenomena. By applying the eigen microstates method20, 21, 22, 23, 24, 25, fundamental patterns and their evolutionary rules can be extracted from system evolution data. In present work, PPS as a class of emergent behaviors in self-organized multi-body systems are studied from a statistical physics perspective, using eigen microstates method to uncover the underlying mechanisms.
@@
The service system can be considered as a general model for many economic and social activities of human beings. In these systems the scheduling of supply is optimized to meet the spatiotemporal characteristics of demand and PPS phenomena is a common feature of these systems. Furthermore, global operational efficiency, cost-effectiveness and preventing system collapse are essential considerations from a societal management point of view. Examples of service systems include transportation systems, like subway, taxis26, buses27, 28 etc; elevators in skyscrapers29, 30, 31, 32; medical appointment systems; logistical management33, 34; aircraft carriers decks management35, 36, 37. In present work, the subway system as a reprehensive example of service system is studied and the mechanism of PPS phenomena of service system is revealed. The results and methodologies may provide valuable insights for service systems.
@@
Figure 1: Stranding phenomena is a major problem of subway system: a The temporal variation in the number of stations with stranded passengers in the subway network throughout a day on May 10, 2021. b Spatial distribution of stations with stranded passengers in the Beijing subway network at 8:10 AM rush hour. c Mechanisms of passengers being stranded at subway stations. The causes for the occurrence of stranded passengers and PPS phenomena are as follows: network topology, service trains scheduling, and spatiotemporal characteristics of demand. d the scheme of causing of PPS phenomena. stranding occurs at stations when passengers’ demand exceeds transportation supply. Due to the interconnected nature of the line, operation in full capacity in upstream stations would occupy the supply scale of downstream stations, leading to passengers stranded at the latter.
2 Results
@@
To investigate the phenomenon of PPS in subway systems, a dynamic model based on an actual subway network is developed, using Beijing as a case study. Beijing is a world-class megacity with a population of 23 million permanent residents. As of July 26, 2022, the Beijing subway network is composed of 17 lines with a total of 330 operational stations and 69 transfer stations, constituting a transportation system of significant scale. The job and residence centers in Beijing exhibit characteristic features of TOD, with a substantial portion of the population relying on the subway for their daily commuting demands. The model system used in the present work, as shown in Fig. 1b, includes 3 east-west linear lines (Lines 1, 6, and 7), 3 north-south linear lines (Lines 4, 5, and 8), and 3 circular lines (Lines 2, 10, and 13), resulting in a subway network comprising 234 stations ($N=234$). Among these stations, there are 36 transfer stations and 10 terminal stations (ends), capturing the representative topological structure of the Beijing subway system. These 9 earliest established lines, operating in the core area of Beijing serve as the main arteries in the real subway system. This model system is an appropriate representation in terms of scale, significance, and topological structure, which enables us to capture the daily operational features. In order to concentrate on the primary physical properties and alleviate computational burden, we have excluded the newly constructed lines from our study that have not yet achieved supply-demand equilibrium.
@@
A dynamic evolutionary model system is proposed within this subway model to quantitatively investigate the phenomenon of PPS. This model system comprises 2 types of active basic units, passengers ($N_{\mathrm{p}}=3,860,008$) and trains ($N_{\mathrm{ori}}=5,266$) with service relationship. The service process includes the operation of trains to meet the demand of passengers. Trains operate according to a predetermined schedule, ensuring punctuality. When a train arrives at a station, passengers on board make decisions to disembark or not according to their destination demand, while passengers waiting at the station decide whether to get onboard based on the crowdedness of the compartment. Through computer simulations, the dynamic trajectories of passengers and trains can be obtained (further details can be found in Supplementary Information, A.2). To elucidate the fundamental physical characteristics of PPS, 2 simulation systems are considered: an ideal model (System I) and a realistic model (System II). In System I, it is assumed that the rates of demand and supply are independent of time and position, and the positions of OD pairs are randomly distributed throughout the network. This model system allows us to study PPS under non-equilibrium steady-state conditions. In System IIII, the rate of passenger demand depends on time, exhibiting with rush hour patterns. The positions of OD pairs capture the tidal nature of real commuting scenarios. The simulation time is measured in seconds, with the system state recorded every 5 minutes for System I and every 10 minutes for System II.
@@
This study employs the average waiting time of passengers and the scale of stranding as the order parameters characterizing the PPS phenomena. The waiting time of passengers, $t_{\mathrm{w}}$, will increase due to the PPS phenomena, which serves as a crucial metric for evaluating the system’s service efficiency. To quantify the PPS phenomena in the subway network, we develop here a quantitative framework based on a physical concept of cluster, which combines evolving traffic dynamics with network structure. Instead of the commonly used structural topology, only stations in the network with stranded passengers are considered functionally connected. In this way, we can characterize and understand the dynamics of PPS phenomena through cluster formation and vanishing. The scale of stranding in the system is characterized by 3 order parameters: the total number of stations within the stranding clusters, $n_{\mathrm{s}}$; the size of the largest cluster, $m_{\mathrm{c}}$, represented by the number of stations within it; and the total number of stranded stations in the system, $n_{\mathrm{t}}$ (quantitative definition of the cluster can be found in Supplementary Information, B.1).